OpenStudy (sneville):

What is the solution of the system? Use elimination. 3x – 4y = –24 x + y = –1 A. (3, 4) B. (4, 3) C. (0, 6) D. (–4, 3)

OpenStudy (australopithecus):

1. solve for one of the variables in equation 1 (pick the easy one) x + y = -1 x = -y - 1 2. Plug that into equation 2 3x – 4y = –24 becomes, 3(-y - 1) – 4y = –24 solve for y 3. Plug the value you got for y into equation 1 or 2 to solve for x